Hi here's the equation (x^3-7x^2+4x+12); (x-2) use subsitution to tell whether the given binomial is a factor of the polynomial.
I keep getting just x^2 then I don't know what to do

Answer by Alan3354(30993)

(x^3-7x^2+4x+12); (x-2) use subsitution to tell whether the given binomial is a factor of the polynomial.
-----------------
Sub 2 for x. If f(2)=0, then x-2 is a factor
f(x) = x^3-7x^2+4x+12
f(2) = 2^3 - 7*2^2 + 4*2 + 12
= 8 - 28 + 8 + 12
= 0
-------
x-2 is a factor.